A train left mooseport traveling west at 50mph.One hour later , another train left Mooseport traveling east at 70 mph. How many hours had the first train been traveling when they were 350 miles apart?
Answer by checkley77(12844)

50t+70(t-1)=350
50t+70t-70=350
120t=350+70
120t=420
t=420/120
t=3.5 hours they will be 350 miles apart.
Proof:
50*3.5+70*2.5=350
175+175=350
350=350